Water is supplied to an axial flow turbine under a Total head of 35 m. The mean diameter of the runner is 2 m and is rotates at 145 rev/metre. Water leaves the guide vanes at 30 degrees to the direction of runner rotation at it mean radius and angle of the runner blade at outlet is 28 degrees. If 7% of the total head is lost in the casing and guide vanes and the relative velocity is reduced by 8% due to friction in the runner, determine the blade angle at inlet (at mean radius) and the hydraulic efficiency of the turbine.
1
Expert's answer
2020-08-11T12:50:44-0400

Given;

H = 35 m , D = 2 m N = 145 rev/ min, alpha = 300 , beta = 280 , hg =7% of H

Head available H = 0.93 * 35 = 32.6 m

inlet velocity V1 = square root of 2gH = sqrt 2*9.8*32.6 = 25.3 m/s

u = (3.14*2*145)/60 = 15.2 m/s

From inlet triangle of velocity

V2r1 = 25.32 + 15.22- (2*15.2*25.3 cos 30) = 14.3 m/s

again

15.22 = 14.32 + 25.32 - ( 2*14.3*25.3 cos (3.14-alpha)) = 1480

Therefore blade angle is 1480

Hydraulic efficiency

C1 = 0.92 sqrt 2gH = 0.92 sqrt 2*9.81*32.6 = 23.26m/s

Cw1 = C1 cos alpha = 23.26 cos 30 = 20.15 m/s

nh = Cw1U/gH = (20.15*15.2)/9.81*32.6 = 0.9577 or 95.8%

Therefore efficiency = 95.8%